# Zeroth-order Informed Fine-Tuning for Diffusion Model: A Recursive Likelihood Ratio Optimizer

### v2

@@ A.7 Proofs
Since the sum of unbiased estimators is still unbiased, it is easy to generalize the result for any number of terms along the lines of the above proof. Also, the unbiasedness of the estimators does not change no matter how we pick the combination of the HO estimator and the ZO estimator. Therefore, we can reach the conclusion that
$\nabla_{\theta}\mathbb{E}[R(\phi_{1:T}(\bm{x}_{T};\theta))]$
$=\mathbb{E}_{\bm{z}\sim f(\bm{z}),j\sim\mathcal{U}(1,T-h)}\bigg{[}\frac{\partial\varphi(\bm{x}_{1};\theta)}{\partial\theta}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x}_{0}}+\sum_{i\in C}R(\bm{x}_{0})\nabla_{\theta}\ln f(\bm{z}_{i})+R(\bm{x}_{0})D^{\top}_{\theta}\phi_{j:j+h}(\bm{x}_{j+h};\theta)\nabla_{\theta}\ln f(\bm{z}_{j})\bigg{]}.$
A.7.4 Variance of the RLR estimator
@@
In this section we discuss the variance of the RLR estimator, which consists of 3 terms, which are denoted as A, B and C respectively:
$RLR=\underbrace{\frac{\partial\varphi(\bm{x}_{1};\theta)}{\partial\theta}}_{A}+$ $\underbrace{R(\bm{x}_{0})D^{\top}_{\theta}\phi_{j:j+h}(\bm{x}_{j+h};\theta)\nabla\ln f(\bm{z}_{j})}_{B}+\underbrace{\sum_{i\in C}R(\bm{x}_{0})\nabla\ln f(\bm{z}_{i})}_{C}.$ (18)
@@
which gives a upper bound for the variance of the RLR estimator. Next we derive the specific expression of the terms $\text{Var}(A)$, $\text{Var}(B)$ and $\text{Var}(C)$.
@@
[…]
@@
By the Hammersley-Chapman-Robbins bound, we can derive an upper bound for the variance:
@@
$\frac{\gamma}{2}\mathbb{E}[\nabla R(\theta_{k})^{2}]\leq\frac{2}{\gamma}(\mathbb{E}(\theta_{k})-\mathbb{E}(\theta_{k+1}))+\gamma L\sigma^{2}_{\mathrm{RLR}}.$
@@
$\frac{1}{K+1}\sum_{k=0}^{K}[\mathbb{E}[\nabla R(\theta_{k})^{2}]]\leq\frac{2(R(x_{0})-R^{\ast})}{K+1}+\gamma L\sigma^{2}_{\mathrm{RLR}}.$
Defining $\Delta_{0}:=f(x_{0})-f^{\ast}$, if we set the step size

### v3

@@ Appendix I More Results on Vbench
Since the sum of unbiased estimators is still unbiased, it is easy to generalize the result for any number of terms along the lines of the above proof. Also, the unbiasedness of the estimators does not change no matter how we pick the combination of the HO estimator and the ZO estimator. Therefore, we can reach the conclusion that:
$\nabla_{\theta}\mathbb{E}_{\bm{z}_{1:3}}[R(\bm{x}_{0})]$ $={D^{\top}_{\theta}\phi(x_{1},z_{1};\theta)}\frac{dR(x_{0})}{dx_{0}}$
$-R(x_{0}){D^{\top}_{\theta}\phi_{j:j+h}(x_{j+h},z_{j:j+h};\theta)}\nabla_{z}\ln f(z_{j})-\sum_{i\in C}R(x_{0})\nabla_{z}\ln f(z_{i}),$
which completes the proof. ∎
J.4 Variance of the RLR Estimator
@@
In this section, we discuss the variance of the RLR estimator, which consists of 3 terms, which are denoted as $A$, $B$, and $C$, respectively:
$RLR=\underbrace{\frac{\partial\phi_{1}(x_{1},z_{1};\theta)}{\partial\theta}}_{A}$- $\underbrace{R(x_{0})D^{\top}_{\theta}\phi_{j:j+h}(x_{j+h},z_{j:j+h};\theta)\nabla\ln f(z_{j})}_{B}-\underbrace{\sum_{i\in C}R(x_{0})\nabla\ln f(z_{i})}_{C}.$ (13)
@@
which gives an upper bound for the variance of the RLR estimator. Next we derive the specific expression of the terms $\text{Var}(A)$, $\text{Var}(B)$ and $\text{Var}(C)$.
@@
$=\lim_{h\to 0}\int\frac{f(z;\theta+h)-f(z;\theta)}{h}R(z)dz$
$=\lim_{h\to 0}\frac{1}{h}\int f(z;\theta)\bigg(\frac{f(z;\theta+h)}{f(z;\theta)}-1\bigg)R(z)dz$
@@
By the Hammersley-Chapman-Robbins bound, we can derive a lower bound for the variance:
@@
$\mathbb{E}[\nabla R(\theta_{k})^{2}]\leq\frac{2}{\gamma}(\mathbb{E}(R(\theta_{k}))-\mathbb{E}(R(\theta_{k+1})))+\gamma L\sigma^{2}_{\mathrm{RLR}}.$
@@
$\frac{1}{K+1}\sum_{k=0}^{K}[\mathbb{E}[\nabla R(\theta_{k})^{2}]]\leq\frac{2(R(x_{0})-R^{\ast})}{\gamma(K+1)}+\gamma L\sigma^{2}_{\mathrm{RLR}}.$
Defining $\Delta_{0}:=R(x_{0})-R^{\ast}$, if we set the step size
@@
J.6 Solution of Optimization Problem (7)
Quadratic Upper Bound on Variance.
Using the variance decomposition bound (14), the objective of the problem (7) can be upper bounded by a quadratic function of $h$:
$Q(h)=a(h+1)^{2}+b(h+1)+c,$ (18)
where the coefficients $a,b,c$ are given by:
$a$ $=V_{h}^{2}+V_{z}^{2}-2{V_{h}V_{z}}>0,$
$b$ $=-2TV^{2}_{z}+2T{V_{h}V_{z}},$
Then, the unconstrained optimal solution $h^{\star}$ is given by:
$h_{\mathrm{unc}}^{\star}=\frac{T(V_{z}^{2}-{V_{h}V_{z}})}{2(V_{h}^{2}+V^{2}_{z}-2{V_{h}V_{z}})}-1.$ (19)
Where $h$ and (T1h)(T-1-h) are the number of steps for HO and ZO; $V^{2}_{h}$ and $V^{2}_{z}$ are coefficients indicating the magnitude of the variance of HO and ZO per step, and $V_{h}\ll V_{z}$ since HO has lower variance. Notice that $T>2$. Therefore, $_{\mathrm{unc}}^{\ast}>0. h\mathcal{B}_{h}$ and $\mathcal{B}_{z}$ are coefficients indicating the magnitude of the memory cost of HO and ZO per step and $\mathcal{B}_{h}>\mathcal{B}_{z}$. $\mathcal{B}_{z}T\leq\mathcal{B}\leq\mathcal{B}_{h}T.$ The constraint of the problem (7) implies that $h\leq\frac{\mathcal{B}-\mathcal{B}_{z}(T-1)}{\mathcal{B}_{h}-\mathcal{B}_{z}},$ and $\frac{\mathcal{B}-\mathcal{B}_{z}(T-1)}{\mathcal{B}_{h}-\mathcal{B}_{z}}>0$.
Then, the solution to the optimization problem is given by:
$\min\{\lfloor\frac{\mathcal{B}-\mathcal{B}_{z}(T-1)}{\mathcal{B}_{h}-\mathcal{B}_{z}}\rfloor,\lfloor\frac{T(V^{2}_{z}-{V_{h}V_{z}})}{2(V_{h}^{2}+V_{z}^{2}-2{V_{h}V_{z}})}-1\rfloor\}.$